The pulmonary forced vital capacity for men between the ages of 20 and 30 has been determined to

        be normally distributed with a mean of 3.8 liters with a standard deviation of 0.7 liters.  Estimate the

        probability that a randomly selected man of age 28 would have a forced vital capacity of 5 liters or

        more. (6 points)
Answer #1

Solution :

Given that,

mean = = 3.8

standard deviation = = 0.7

P(x >5 ) = 1 - P(x<5 )

= 1 - P[(x -) / < (5-3.8) / 0.7]

= 1 - P(z < 1.71)

Using z table

= 1 -  0.9564

probability= 0.0436